# Understanding differences in applying DETR to natural and medical images

### v1

@@ Abstract
Transformer-based detectors have shown success in computer vision tasks with natural images. These models, exemplified by the Deformable DETR, are optimized through complex engineering strategies tailored to the typical characteristics of natural scenes. However, medical imaging data presents unique challenges such as extremely large image sizes, fewer and smaller regions of interest, and object classes which can be differentiated only through subtle differences. This study evaluates the applicability of these transformer-based design choices when applied to a screening mammography dataset that represents these distinct medical imaging data characteristics. Our analysis reveals that common design choices from the natural image domain, such as complex encoder architectures, multi-scale feature fusion, query initialization, and iterative bounding box refinement, do not improve and sometimes even impair object detection performance in medical imaging. In contrast, simpler and shallower architectures often achieve equal or superior results. This finding suggests that the adaptation of transformer models for medical imaging data requires a reevaluation of standard practices, potentially leading to more efficient and specialized frameworks for medical diagnosis.
Keywords: Machine Learning, Vision Transformers, Object Detection, Breast Cancer
@@ 1 Introduction
Figure 1: An overview of the study. In this work, we examine design choices in Deformable DETR. To evaluate them, we used the NYU Breast Cancer Screening Dataset. Specifically, we assess six design choices (highlighted in red): 1. input resolution; 2. number of encoder layers; 3. whether to use multi-layer feature fusion; 4. number of object queries; 5. query initialization method; 6. whether to use the iterative bounding box refinement method. The graph on the right demonstrates the performance changes, indicated by Average Precision (AP), after applying these design choices. +(2): reducing the number of encoder layers; +(3): removing multi-layer feature fusion; +(4): reducing the number of object queries; +(5): adding object query initialization method; +(6): adding iterative bounding box refinement.) The results reveal that a simplified architecture +(2,3,4) is more suitable for medical datasets and results in improved performance. Increasing complexity of the model, as indicated by +(2,3,4,5,6), brings marginal benefit.
@@
Recent advancements in computer vision have increasingly focused on leveraging transformer architectures (Vaswani et al., 2017) for image classification and object detection (Dosovitskiy et al., 2020; Liu et al., 2021; Carion et al., 2020; Touvron et al., 2021). Thanks to their inherent self-attention mechanisms, transformers effectively capture global dependencies and understand contextual relations across entire images, enabling them to acquire more expressive features. Owing to these advantages, transformer-based models have become prevalent in natural imaging analysis. Moreover, their application in medical imaging has shown promising results, indicating their potential to transform this field as well. (Chen et al., 2021; Dai et al., 2021b; Valanarasu et al., 2021; Zheng et al., 2022).
Object detection plays a crucial role in medical image analysis as detection models directly identify the locations of abnormalities, which is critically important in application of AI in medical diagnosis. With the rise of transformer-based models, Detection Transformer (DETR), an end-to-end transformer-based object detection framework (Carion et al., 2020) has become popular. DETR does not require Non-Maximum Suppression (NMS) (Girshick et al., 2014), a post-processing step that reduces overlapping detections, which is non-differentiable and sensitive to predefined thresholds. The elimination of NMS enables end-to-end training and direct optimization of the objective function. DETR has established itself as one of the most competitive detection framework by consistently achieving state-of-the-art accuracy on MS-COCO datasets (Zhu et al., 2020; Zhang et al., 2022; Zong et al., 2023). Consequently, the potential of DETR draw significant research interest in improving overall performance and accelerating learning (Zhu et al., 2020; Zhang et al., 2022; Chen et al., 2022b; Wang et al., 2022; Chen et al., 2022a). This has led to the development of some highly engineered DETR-based architectures and a diverse range of design choices within the DETR family.
Despite the success of DETR architectures for natural image data, their application to medical imaging poses challenges. Medical images differ significantly from natural images in several aspects:
  1. 1.
@@
    High resolution and small regions of interest Medical images are usually in very high resolutions with small, critical regions of interest (e.g., lesions, masses and calcifications) (Moawad et al., 2023; Heath et al., 2001).
  2. 2.
    Standardized acquisition Medical images are acquired through very standardized procedures, resulting in images consisting of three elements: the body part of interest, the surrounding body tissue, and black pixels outside the body. In contrast, the backgrounds in natural images can contain anything existing in the natural world.
  3. 3.
    Small number of objects per image Medical images typically focus on a narrow range of abnormalities, resulting in fewer objects of interest within each image from very few classes.
  4. 4.
@@
    Small and imbalanced data sets Medical imaging data sets are often smaller and exhibit a more imbalanced class distribution, as positive cases (i.e. unhealthy subjects) are usually a lot less common than negative cases (i.e. healthy subjects)(Galdran et al., 2021; Heath et al., 2001; Wang et al., 2017).
Many highly engineered design choices of DETR-family models, such as multi-scale feature fusion, iterative bounding box refinement, works well for natural image data. However, when applying to medical images, where data characteristics differ markedly, we need to reexamine the effectiveness of those engineering design choices. Based on the differences between medical and natural images, we hypothesize that those design choices make DETR architecture unnecessarily complex, increasing its computation demand and memory consumption, and might not even work well with the unique properties of medical imaging data listed above. We hypothesize that a simplified model, tailored to the characteristics of medical imaging data, might be more efficient and accurate.
To verify the aforementioned hypothesis, this work investigates whether design choices in the DETR family, originally intended for natural images, remain beneficial or even necessary for medical imaging applications. We use the Deformable DETR as our baseline model and the NYU Breast Cancer Screening Dataset as a representative case study (Wu et al., 2019). This dataset contains high-resolution screening mammography with small lesions, serving as a representative dataset for the characteristics of medical imaging mentioned above. The key findings of our work are:
@@
  • Models with a reduced number of encoder layers and no multi-scale feature fusion learn faster without compromising detection performance. These changes actually increased performance by $1.1\%$ in $\mathrm{AP}_{10,50}$, while accelerating the training by 50%.
@@
  • Additional techniques applied in the decoder, such as object query initialization methods and iterative bounding box refinement, do not improve detection performance on medical image datasets as they do on natural image datasets. Instead, they resulted in a 0.7% drop in $\mathrm{AP}_{10,50}$, indicating these techniques may not be suitable for medical imaging tasks.
  • Increasing the number of object queries improves detection performance up to a point. However, unlike for natural image data, using an excessive number of queries (e.g., 800) leads to a 2% drop in performance due to an increase in false positives.
  • Design choices aimed at localizing ROI candidates, such as object query initialization methods, iterative bounding box refinement and more object queries, do not improve performance, suggesting that correctly classifying ROI candidates is key to performance improvement.
@@ 2 Background on DETRs
DETR offers several advantages over commonly applied detection models such as Mask-RCNN (He et al., 2017) and YOLO (Redmon et al., 2016). Its transformer-based architecture enables learning more expressive features than CNN-based models, and end-to-end training improves optimization and performance. However, DETR suffers from slow learning. To address this issue, several studies have proposed techniques to accelerate learning and enhance overall performance (Zhu et al., 2020; Wang et al., 2022; Chen et al., 2022b; Zhang et al., 2022). Deformable DETR stands out for its competitive performance on the MS COCO dataset (Lin et al., 2014). It introduces a deformable attention module, which reduces training time by a factor of 10 and enables multi-scale feature fusion that improve detection, especially for small objects. Given its widespread adoption as a benchmark in the subsequent research works (Roh et al., 2021; Dai et al., 2021a; Zhang et al., 2022; Yao et al., 2021), we choose Deformable DETR as the baseline architecture for our experiments. This section details the essential components of DETR and Deformable DETR architectures.
@@
DETR consists of a backbone, an encoder-decoder transformer, and a prediction head, as illustrated in Figure 2(a).
Given an input image $x\in\mathbb{R}^{C_{0}\times W_{0}\times H_{0}}$, the backbone network $f$ produces a low-resolution activation map $x_{s}=f(x)\in\mathbb{R}^{C\times W\times H}$. This map is further processed by a $1\times 1$ convolution to collapse the channel dimension $C$ into a smaller size $d$, resulting in image tokens $x_{f}=\mathrm{conv}(x_{s})\in\mathbb{R}^{WH\times d}$. To preserve spatial information in the original image, each token is paired with a positional encoding, denoted by $x_{p}\in\mathbb{R}^{WH\times d}$. The encoder is a standard attention-based transformer where each layer consists of a multi-head self-attention module (MHSA) followed by a feedforward network (FFN). For an in-depth formalization of MHSA, refer to Appendix A.1. Typically, the DETR encoder consists of 6 layers. The encoder preserves the dimension of the input, producing $x_{enc}\in\mathbb{R}^{WH\times d}$.
The decoder receives two inputs, the encoded features $x_{enc}$ and $N$ object queries $q\in\mathbb{R}^{N\times d}$. Object queries play a central role in DETR architecture. They are learnable embeddings that work as placeholders for the potential objects in an image. Each of them attends to the specific regions of the image and is individually decoded into a bounding box prediction. Each object query is the sum of two learnable embedding: content embeddings $q_{c}\in\mathbb{R}^{N\times d}$, initialized as zero vectors, and the positional embeddings $q_{p}\in\mathbb{R}^{N\times d}$, indicating each query’s position. More methods for initializing object queries are discussed in Section 3. Decoder layers consists of a MHSA, enabling inter-query learning, and multi-head (MH) cross-attention to integrate encoder features, and a FFN. The formalization of MH cross-attention is detailed in Appendix A.2.
@@
Figure 2: Architecture of DETR and Deformable DETR. Both architectures consist of a backbone, an encoder-decoder transformer and a prediction head. Deformable DETR differs from DETR in its utilization of multiple feature level fusions and its application of deformable attention instead of standard attention mechanism.
@@
Deformable DETR improves upon DETR by introducing a deformable attention module, which accelerates training and enhances the detection of small objects. The architecture of Deformable DETR is illustrated in Figure 2(b).
@@
For dense prediction tasks such as object detection, incorporating higher-resolution feature maps can substantially improve detection performance, especially for smaller objects (He et al., 2017). However, the complexity of the standard attention mechanism is quadratic with respect to the number of tokens, making it infeasible for multiple scales of feature maps. The deformable attention mechanism enables effective multi-scale feature fusion. Specifically, the encoder receives the output feature maps $x_{1},x_{2},x_{3}$ from the backbone’s last three layers, and a convolutional layer generates the lowest resolution feature map $x_{4}$. All four feature maps undergo a $1\times 1$ convolution and then are reshaped into a sequence of feature vectors of dimension $d$, denoted by $x_{f}\in\mathbb{R}^{M\times d}$. Each token is associated with a positional embedding, as well as a layer embedding to identify feature map level. Section 3 explores the benefits of multi-feature fusion for medical imaging datasets.
Moreover, Deformable DETR introduces reference points in the deformable attention module. In the encoder, each query $q$ is associated with a 2D reference point $p_{q}=[x,y]$, denoting its location on the feature map. The key sampling function generates $k$ sampling offsets with respect to the reference point, and thus determines the $k$ keys for the query. Similarly in the decoder, the reference point of each object query $q$ is defined by a linear projection of its positional embedding $q_{p}$. In this way, each object query can be mapped to a position on the feature map. This approach allows object queries to focus on specific regions, significantly accelerating learning Zhu et al. (2020).
@@
DETR-based architectures have been widely applied to various medical imaging tasks, often with architectural tweaks to improve overall performance. For example, Mathai et al. (2022) leveraged a bounding box fusion technique on DETR to reduce false-positive rate in lymph nodes detection. MyopiaDETR (Li et al., 2023) utilizes a Feature Pyramid Network to improve detection of small objects in lesion detection of pathological myopia. COTR (Shen et al., 2021) embeds convolutional layers into DETR encoders to accelerate learning in polyp detection. Although these works achieved good performances, our experiments indicates that, contrary to the common understanding, simplifying the DETR architecture can improve accuracy and accelerate training. One work we identified that also points in this direction, Cell-DETR (Prangemeier et al., 2020), also reduces the number of parameters by ten times, achieving faster inference speeds while maintaining performance on par with state-of-the-art baselines. Finally, Garrucho et al. (2023) applied out-of-the-box Deformable DETR on mammography for mass detection. However, their focus is the effect of a data augmentation method on its detection performance. Despite these advances, a systematic exploration of the effectiveness and relevance of foundational DETR design choices remains under-explored.

### v2

@@ Abstract
Natural images depict real-world scenes such as landscapes, animals, and everyday items. Transformer-based detectors, such as the Detection Transformer, have demonstrated strong object detection performance on natural image datasets. These models are typically optimized through complex engineering strategies tailored to the characteristics of natural scenes. However, medical imaging presents unique challenges, such as high resolutions, smaller and fewer regions of interest, and subtle inter-class differences, which differ significantly from natural images. In this study, we evaluated the effectiveness of common design choices in transformer-based detectors when applied to medical imaging. Using two representative datasets, a mammography dataset and a chest CT dataset, we showed that common design choices proposed for natural images, including complex encoder architectures, multi-scale feature fusion, query initialization, and iterative bounding box refinement, fail to improve and can even be detrimental to the object detection performance. In contrast, simpler and shallower architectures often achieve equal or superior results with less computational cost. These findings highlight that standard design practices need to be reconsidered when adapting transformer models to medical imaging, and suggest that simplicity may be more effective than added complexity in this domain. Our model code and weights are publicly available at https://github.com/nyukat/Mammo-DETR.
doi:
https://doi.org/10.59275/j.melba.2025-g137
keywords:
Deep Learning, Machine Learning, Vision Transformers, Object Detection, Breast Cancer
volume: 3
@@ 1 Introduction
Figure 1: An overview of the study. In this work, we investigate key design choices in Deformable DETR using the NYU Breast Cancer Screening Dataset (NYU Breast) and LUNA16 Dataset. Specifically, we evaluate six design factors (highlighted in red): (1) input resolution, (2) number of encoder layers, (3) use of multi-layer feature fusion, (4) number of object queries, (5) query initialization method, and (6) use of iterative bounding box refinement. The graph on the right shows changes in Average Precision (AP) resulting from these design choices. +2 indicates a reduction in encoder layers; +3 removes multi-layer fusion; +4 reduces object queries; +5 adds query initialization; and +6 enables iterative box refinement. Our findings suggest that a simplified architecture (+2,3,4) is better suited for medical datasets, leading to improved performance.
@@
Recent advances in computer vision have increasingly turned to transformer architectures (Vaswani et al., 2017) for tasks such as image classification and object detection (Dosovitskiy et al., 2020; Liu et al., 2021; Carion et al., 2020; Touvron et al., 2021). With their inherent self-attention mechanisms, transformers effectively capture global dependencies and understand contextual relations across the entire image. These strengths have made transformer-based models a popular choice in natural image analysis. Their application in medical imaging has shown promising results, suggesting strong potential in this field as well (Chen et al., 2021; Dai et al., 2021b; Valanarasu et al., 2021; Zheng et al., 2022).
Object detection is crucial in medical image analysis, as detection models identify the locations of abnormalities, which are important for medical diagnosis. Among transformer-based detectors, Detection Transformer (DETR) (Carion et al., 2020) has gained popularity for its end-to-end training pipeline and elimination of non-differentiable post-processing steps such as Non-Maximum Suppression (NMS) (Girshick et al., 2014). By leveraging the transformer architecture and directly optimizing the objective function, DETR achieves state-of-the-art results on natural image benchmarks such as MS COCO (Zhu et al., 2020; Zhang et al., 2022; Zong et al., 2023). Its success has drawn intense research interest, leading to a range of highly engineered DETR variants aimed at boosting accuracy and training efficiency (Zhu et al., 2020; Zhang et al., 2022; Chen et al., 2022b; Wang et al., 2022; Chen et al., 2022a).
Despite the success of DETR architectures on natural image benchmarks, their direct application to medical imaging remains challenging due to fundamental differences between the two domains (Figure 2):
@@
  • High resolution and small regions of interest: Medical images are often extremely high-resolution, with clinically relevant features, such as lesions or calcifications, occupying only small portions of the image (Moawad et al., 2023; Heath et al., 2001).
  • Standardized acquisition protocols: Unlike natural images which have diverse backgrounds, medical images are acquired under standardized procedures, resulting in consistent anatomical structures and minimal background variability.
  • Few objects per image: Medical images usually focus on a narrow range of abnormalities, resulting in fewer objects of interest and a narrower class space compared to the rich and diverse class space of natural images. Additionally, many medical images may not contain any objects at all.
@@
  • Small and imbalanced data sets: Medical imaging data sets are often small and exhibit a more unbalanced class distribution, as positive cases (i.e., unhealthy subjects) are usually much less common than negative cases (i.e., healthy subjects)(Galdran et al., 2021; Heath et al., 2001; Wang et al., 2017).
DETR-family models, such as Deformable DETR (Zhu et al., 2020), incorporate complex design choices such as multi-scale feature fusion and iterative bounding box refinement to address challenges in natural image detection. However, their effectiveness in medical imaging is unclear, as the domain presents distinct characteristics: high resolution, small lesion size, limited object diversity, and class imbalance, which differ markedly from natural images. In such settings, detecting subtle features precisely is often more important than modeling diverse object scales or dense scenes. As a result, these complex design choices may introduce unnecessary computational overhead and memory cost without yielding performance gains.
In this study, we examine how DETR can be adapted to better suit medical imaging tasks. We hypothesize that a simplified model, tailored to the specific characteristics of medical data, can achieve comparable performance with reduced computational cost. To evaluate this, we use Deformable DETR (Zhu et al., 2020) as a baseline on two medical imaging datasets: the NYU Breast Cancer Screening Dataset (Wu et al., 2019) and the LUNA16 dataset (Setio et al., 2017), a public chest CT dataset focused on lung nodule detection (Figure 2). Those two datasets highlight the distinct features of medical images, such as high resolution, small lesions, and class imbalance.
Our experiments demonstrate that simplified DETR configurations—using fewer encoder layers, a single feature map, and no decoding enhancements—achieve detection performance on par with, or better than, standard Deformable DETR, while substantially reducing computational cost. These findings validate our hypothesis and highlight the potential of lightweight DETR variants as efficient and effective baselines for medical imaging. The key findings of our work are:
@@
  • Models with a reduced number of encoder layers and no multi-scale feature fusion learn faster without compromising detection performance. These changes maintain performance within $1\%$ in $\mathrm{AP}_{10,50}$ on both datasets, while accelerating training by up to 40%.
  • Increasing the number of object queries to around 100 queries improves localization and detection performance. Beyond this point, performance declines, primarily due to a rise in false positives that obscure true positive detections.
@@
  • Decoding techniques such as object query initialization and iterative bounding box refinement, while beneficial for natural image detection, do not improve performance on medical datasets. In some cases, they degrade performance (e.g., a 0.7% drop in $\mathrm{AP}_{10,50}$ for NYU Breast and 1.8% drop for LUNA16), likely due to overfitting and limited positive examples.
Refer to captionRefer to caption
Figure 2: Example images from the LUNA16, NYU Breast Cancer Screening, and MS COCO datasets. (a) A chest CT slice from LUNA16 showing a lung nodule (red box). Among images that contain nodules, the average number of objects per image is 1.21. This dataset has one class. (b) A mammogram from the NYU Breast Cancer Screening dataset showing a cancerous lesion (red highlight). In images containing lesions, there are on average 1.10 objects per image. This dataset has two classes. (c) Two example images from the MS COCO dataset. These illustrate the complexity of natural scenes, with multiple overlapping objects of varying sizes. On average, MS COCO images contain 7.33 objects per image. MS COCO has 80 object classes.
@@ 2 Background on DETRs
DETR (Carion et al., 2020) offers several advantages over traditional detection models such as Mask-RCNN (He et al., 2017) and YOLO (Redmon et al., 2016). Its transformer-based architecture enables more expressive feature representations, and its end-to-end training simplifies optimization and improves performance. However, DETR suffers from slow learning. To address this issue, various extensions have been proposed to accelerate training and improve detection performance (Zhu et al., 2020; Wang et al., 2022; Chen et al., 2022b; Zhang et al., 2022). Deformable DETR (Zhu et al., 2020) stands out for its competitive performance on the MS COCO dataset (Lin et al., 2014). It introduces a deformable attention module that reduces training time by a factor of 10 and enables multi-scale feature fusion that improves detection, especially for small objects. Given its strong performance and widespread adoption in subsequent research (Roh et al., 2021; Dai et al., 2021a; Zhang et al., 2022; Yao et al., 2021), we adopt Deformable DETR as the baseline for our experiments. This section outlines the key components of DETR and Deformable DETR architectures.
@@
DETR (Carion et al., 2020) consists of a backbone, an encoder-decoder transformer, and a prediction head, as illustrated in Figure 3(a).
Given an input image $x\in\mathbb{R}^{C_{0}\times W_{0}\times H_{0}}$, where $C_{0}$ is the number of channels and $W_{0}$ and $H_{0}$ are the width and height, the backbone network $f$ produces a low-resolution activation map $x_{s}=f(x)\in\mathbb{R}^{C\times W\times H}$, with $C$ significantly larger than $C_{0}$. The specific sizes of $W,H$ and $C$ depend on the choice of backbones. For instance, when using Swin-T (Liu et al., 2021) as the backbone, the spatial dimensions are downsampled to $W=W_{0}/32$ and $H=H_{0}32$, and the number of channels is increased to $C=768$. This map is further processed by a $1\times 1$ convolution to collapse the channel dimension $C$ into a smaller size $d$, resulting in image tokens $x_{f}=\mathrm{conv}(x_{s})\in\mathbb{R}^{WH\times d}$. To preserve spatial information in the original image, each token is paired with a positional encoding, denoted by $x_{p}\in\mathbb{R}^{WH\times d}$. The encoder is a standard attention-based transformer where each layer consists of a multi-head self-attention module (MHSA) followed by a feedforward network (FFN). For an in-depth formalization of MHSA, refer to the Appendix A.1. Typically, the DETR encoder consists of 6 layers. The encoder preserves the dimension of the input, producing $x_{enc}\in\mathbb{R}^{WH\times d}$.
The decoder receives two inputs, the encoded features $x_{enc}$ and $N$ object queries $q\in\mathbb{R}^{N\times d}$. Object queries play a central role in the DETR architecture. They are learnable embeddings that work as placeholders for potential objects in an image. Each of them attends to the specific regions of the image and is individually decoded into a bounding box prediction. Each object query is the sum of two learnable embeddings: content embeddings $q_{c}\in\mathbb{R}^{N\times d}$, initialized as zero vectors, and positional embeddings $q_{p}\in\mathbb{R}^{N\times d}$, indicating each query’s position. More methods for initializing object queries are discussed in Section 3. Decoder layers consists of a MHSA, enabling inter-query learning, and multi-head (MH) cross-attention to integrate encoder features, and an FFN. The formalization of MH cross-attention is detailed in the Appendix A.2.
@@
Figure 3: Architecture of DETR and Deformable DETR. Both architectures consist of a backbone, an encoder-decoder transformer and a prediction head. Deformable DETR differs from DETR in its utilization of multiple feature level fusions and its application of deformable attention instead of standard attention mechanism. Abbreviations: MHSA: multi-head self-attention module; MH cross attention: multi-head cross-attention; MH Deformable SA: multi-head deformable self-attention module; MH Deformable CA: multi-head deformable cross-attention module; FFN: feedforward network.
@@
Deformable DETR (Zhu et al., 2020) improves upon DETR by introducing a deformable attention module, which accelerates training and enhances the detection of small objects. The architecture of Deformable DETR is illustrated in Figure 3(b).
@@
For dense prediction tasks such as object detection, incorporating higher-resolution feature maps can substantially improve detection performance, especially for smaller objects (He et al., 2017). However, the complexity of the standard attention mechanism is quadratic with respect to the number of tokens, making it infeasible for multiple scales of feature maps. The deformable attention mechanism enables effective multi-scale feature fusion. Specifically, the encoder receives the output feature maps $x_{1},x_{2},x_{3}$ from the last three layers of the backbone, and a convolutional layer generates the lowest resolution feature map $x_{4}$. All four feature maps undergo a $1\times 1$ convolution and then are reshaped into a sequence of feature vectors of dimension $d$, denoted by $x_{f}\in\mathbb{R}^{M\times d}$. Each token is associated with a positional embedding, as well as a layer embedding to identify feature map level. Section 3 explores the benefits of multi-feature fusion for medical imaging datasets.
Moreover, Deformable DETR introduces reference points in the deformable attention module. In the encoder, each query $q$ is associated with a 2D reference point $p_{q}=[x,y]$, denoting its location on the feature map. The key sampling function generates $k$ sampling offsets with respect to the reference point, and thus determines the $k$ keys for the query. Similarly in the decoder, the reference point of each object query $q$ is defined by a linear projection of its positional embedding $q_{p}$. In this way, each object query can be mapped to a position on the feature map. This approach allows object queries to focus on specific regions, significantly accelerating learning (Zhu et al., 2020).
@@
DETR-based architectures have been widely applied to various medical imaging tasks, often with architectural tweaks to improve overall performance. For example, Mathai et al. (2022) leveraged a bounding box fusion technique in DETR to reduce the false positive rate in lymph nodes detection. MyopiaDETR (Li et al., 2023) utilizes a Feature Pyramid Network to improve the detection of small objects in lesion detection of pathological myopia. COTR (Shen et al., 2021) embeds convolutional layers into DETR encoders to accelerate learning in polyp detection. Although these works achieved good performances, our experiments indicate that, contrary to the common understanding, simplifying the DETR architecture can improve accuracy and accelerate training. We identified a work that also points in this direction, Cell-DETR (Prangemeier et al., 2020), also reduces the number of parameters tenfold, achieving faster inference speeds while maintaining performance on par with state-of-the-art baselines. Finally, Garrucho et al. (2023) applied out-of-the-box Deformable DETR on mammography for mass detection. However, their focus is the effect of a data augmentation method on its detection performance. Despite these advances, a systematic exploration of the effectiveness and relevance of foundational DETR design choices remains underexplored.